OpenStudy (anonymous):

In a nuclear reaction, the energy released is equal to 8.1 x 1016 joules. Calculate the mass lost in this reaction. (1 J = 1 kg m2/s2) Answer 7.2 x 1032 kg 2.7 x 108 kg 9.0 x 10-1 kg 7.2 x 10-16 kg

OpenStudy (anonymous):

How do I calculate mass lost?

OpenStudy (anonymous):

you have E= 5.4 x 10^(15) given, and speed of light 'c'= 3 x 10^8 m/sec. So, just plug these in "E=(deltaM)C^2, and calculate 'deltaM', which is the mass lost. P.S. - delta just means 'change' so 'deltaM' means the change in mass nothing else.

OpenStudy (anonymous):

Just look at this http://openstudy.com/study#/updates/4fc72c0ce4b022f1e12eac37

OpenStudy (anonymous):

prob. B, yes?

OpenStudy (anonymous):

please help

OpenStudy (anonymous):

I don't understand how to input delta in an equation...how do I do that? how??

OpenStudy (freethinker):

because you didn't read anything about mass defect

OpenStudy (anonymous):

gaahhhhhh......I don't understand their one paragraph description of the concept.

OpenStudy (anonymous):

alright, it's the mass lost during a reaction that converts to energy

OpenStudy (anonymous):

I think it's B

OpenStudy (jfraser):

do the same thing as your last problem

OpenStudy (freethinker):

yeh binding energy \[E = \Delta mc^2 \] we already know the amount of energy released and we also know the speed of light E = 8.1 x 1016 joules \[E = \Delta M c^2 = \Delta M (2.9987 \times 10^2 mc^{-1})^2\]

OpenStudy (freethinker):

I changed the mass into capital M since I am including units of meter/sec

OpenStudy (anonymous):

ok...

OpenStudy (freethinker):

\[8.1 \times 10^{16} kgm^2s^{-2}= \Delta M (2.9987 \times 10ms^{-1})^2\]

OpenStudy (freethinker):

see how easy it is to solve for the mass defect to be calculated?

OpenStudy (anonymous):

hahaha....yeah, it's a freakin cake walk (puts head between knees)

OpenStudy (freethinker):

I keep messing up the exponents, just fix them for me

OpenStudy (freethinker):

I still have to get used to latex

OpenStudy (freethinker):

make sure you have the right data on energy released

OpenStudy (freethinker):

8.1 x 1016 joules is not correct

OpenStudy (freethinker):

\[8.1 \times 10^{16}kgm^2s^{-2} = \Delta M (2.9979 \times 10^{8} ms^{-1})^2\]

OpenStudy (jfraser):

since \[E = \Delta mc^2\] then \[\Delta m = \frac{E}{c^2}\]and \[\Delta m = \frac{8.1*10^{16}J}{(3.0*10^8\frac{m}{s})^2}\] so \[\Delta m = 0.9kg\]

OpenStudy (anonymous):

I don't...I don't....gah... how do I solve delta? what do I do there? It's not an imput on the calculator, and I'm not aware of any value for it.. oh wait, JF just replied..

OpenStudy (freethinker):

the units does not match @JFraser you need to fix it first

OpenStudy (anonymous):

8.1 x 10^16

OpenStudy (jfraser):

the units do match

OpenStudy (anonymous):

A

OpenStudy (freethinker):

if you fix it in the end it will match, but he needs to see it

OpenStudy (anonymous):

ooohh....intellect battle.. (chews nails excitedly)

OpenStudy (jfraser):

i just did it step by step so he can see it

OpenStudy (freethinker):

okay

OpenStudy (freethinker):

the delta just means "change" it is just a symbol

OpenStudy (anonymous):

so...ignore it?

OpenStudy (jfraser):

no, don't ignore it, understand it

OpenStudy (freethinker):

wel you can't really ignore it because it tells you a great deal

OpenStudy (anonymous):

I think you're too smart for the calculator - WolframAlpha can't understand you, lol.

OpenStudy (jfraser):

you're not solving for the mass of either the reactants or the products, you're solving for the difference between the two. that's why the delta is important

OpenStudy (freethinker):

it tells you that there is a change

OpenStudy (anonymous):

A, I think

OpenStudy (freethinker):

no wonder you're not understanding the concepts, you're relying on wolframalpha

OpenStudy (anonymous):

no, no...maybe

OpenStudy (anonymous):

not entirely tho, it helps with longer equations where I'm more liable to make mistakes

OpenStudy (freethinker):

read the chapter 14 on that whole chemistry book link I gave you it is an intro to nuclear chemistry

OpenStudy (anonymous):

what do the ms and s stand for?

OpenStudy (freethinker):

I just don't understand how are you taking chemistry at this point without knowing what deltas are

OpenStudy (freethinker):

m = meters s = seconds

OpenStudy (anonymous):

ok....I think we're over-complicating this - I don't think the equation they want me to use is this complicated...hold on

OpenStudy (freethinker):

@JFraser is using a fractional format, I am using a one line so you will notice negative exponents in my equations

OpenStudy (freethinker):

THIS IS NOT COMPLICATED! this is Algebra 1

OpenStudy (anonymous):

BWHAHAHAHAAHA!! I GET Algebra...this is just mixing math with little symbols and numbers that you have to constantly flick back and fro to the textbook to understand.

OpenStudy (anonymous):

no...it really wasn't that complicated...I still don't know how to solve this freakin thing, but it most certainly is not the way he was describing it. Overc-omplication - can't judge tho - do it myself.

OpenStudy (anonymous):

*over-complication

OpenStudy (jfraser):

is it the importance of the delta that's confusing you, or the algebraic rearrangements of the pieces?

OpenStudy (anonymous):

I don't understand how to calculate m and s. It's fine to have variable in an equation, but I need to know how to find them And they didn't give an equation teaching me how to calculate m and s from the amount of energy released.

OpenStudy (anonymous):

I finally sent an email to my teacher about it, hopefully she'll be able to help.

OpenStudy (anonymous):

The equation they give me to use is... (1 J = 1 kg m2/s2)

OpenStudy (anonymous):

I know J...but not m or s hence my confusion

OpenStudy (freethinker):

the equivalent unit \[1.0 J = 1.0 \frac{ kgm^2 }{ s^2 }\] (kilograms*meters^2) per seconds^2 this is a unit most students learn from physics

OpenStudy (freethinker):

I honesty think that this level of chemistry is not for you you still need to get your elementary algebra situated

OpenStudy (jfraser):

the "m" and "s" inside the equation you gave aren't values to find, they are part of the unit itself. the unit of energy (J) is made up of smaller units (kg, m, and s) in a particular combination that looks like what you posted above. You really need help recognizing units, too